OpenStudy (anonymous):

How many solutions are there for the following system? y = 3x - 4 y = 2x + 1

OpenStudy (anonymous):

i thnk its 0

OpenStudy (anonymous):

oh thanks!

OpenStudy (anonymous):

It is one solution and it is (5,11) because if you graph it, that is where they intersect :)

OpenStudy (anonymous):

1 solution

OpenStudy (anonymous):

yes:)

OpenStudy (unklerhaukus):

3x - 4 = 2x + 1 (3-2)x = 1+ 4

OpenStudy (anonymous):

the intersection point is (5,11).so it has 1 solution

OpenStudy (denisaboichuk):

0